OpenStudy (anonymous):

if sin a = -3/5 and lies in 3 quadrant then the value of cos(a/2) is?

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aum):



OpenStudy (anonymous):

ok then?

OpenStudy (aum):

Since "a" is in the third quadrant, cosine will also be negative and it is -4/5 from the above diagram. Using the half angle formula: cos(a) = 2cos^2(a/2) - 1. -4/5 = 2cos^2(a/2) - 1 1 - 4/5 = 2cos^2(a/2) 2cos^2(a/2) = 1/5 cos^2(a/2) = 1/10 cos(a/2) = plus or minus square root of 1/10 Since a is in the third quadrant, a/2 will be in the second quadrant where cosine is negative. Therefore, \(\cos(a/2) = -\frac{1}{\sqrt{10}} \)

OpenStudy (anonymous):

okey thanku